OpenStudy (anonymous):

Find the set of values of k for which the line y=2x-5 cuts the curve y=x^2+kx+11 in two distinct points.

sam (.sam.):

b^2-4ac

sam (.sam.):

y=2x-5, y=x^2+kx+11 x^2+kx+11=2x-5 x^2-2x+kx+16=0 x^2+(-2+k)x+16=0 a=1 b=-2+k c=16

sam (.sam.):

b^2-4ac= (-2+k)-4(1)(16)=0 solve

OpenStudy (anonymous):

Thank you very much (:

sam (.sam.):

yw